OpenStudy (anonymous):

OK I have two problems I'm stuck on can someone plz help me out and plz teach me how you solved it If b = 12, find c. If b = 3, find a.

OpenStudy (anonymous):

its a triangle btw

OpenStudy (anonymous):

feels like the question is missing something

OpenStudy (anonymous):

the Picture but i don't think you need it

OpenStudy (anonymous):

yes you need the pic ....its tells you the angles which tells the ratios of the sides to on another

OpenStudy (anonymous):

have you learned SOHCAHTOA? (theta\[\sin \Theta=opp/hyp \] \[\cos \Theta=adj/hyp\] \[\tan \theta=opp/adj\]

OpenStudy (anonymous):